# Towards Anomaly-Aware Pre-Training and Fine-Tuning for Graph Anomaly Detection

### v2

@@ Appendix D Proof of Theorem 1
$\langle\boldsymbol{\hat{X}}_{i},\boldsymbol{w}^{*}\rangle+b^{*}=\left\{\begin{aligned} -\frac{R(p_{2}-q_{2})}{2(p_{2}+q_{2})}\|\boldsymbol{\mu}-\boldsymbol{\nu}\|(1+o(1))>0\quad\text{for}\,i\in\mathcal{C}_{a}\cap\mathcal{H}_{e},\\ -\frac{R(q_{2}-p_{2})}{2(p_{2}+q_{2})}\|\boldsymbol{\mu}-\boldsymbol{\nu}\|(1+o(1))<0\quad\text{for}\,i\in\mathcal{C}_{n}\cap\mathcal{H}_{e}.\\ \end{aligned}\right.$ (20)
Next, consider the high-pass filtered node features $\boldsymbol{\hat{X}}=-\boldsymbol{D}^{-1}\boldsymbol{A}\boldsymbol{X}$, the mean of nodes in different classes and patterns are given by:
$\mathbb{E}(\boldsymbol{\hat{X}}_{i})=\left\{\begin{aligned} -\frac{p_{1}\boldsymbol{\mu}+q_{1}\boldsymbol{\nu}}{p_{1}+q_{1}}(1+o(1))\quad\text{for}\,i\in\mathcal{C}_{a}\cap\mathcal{H}_{o},\\ -\frac{q_{1}\boldsymbol{\mu}+p_{1}\boldsymbol{\nu}}{p_{1}+q_{1}}(1+o(1))\quad\text{for}\,i\in\mathcal{C}_{n}\cap\mathcal{H}_{o},\\ -\frac{p_{2}\boldsymbol{\mu}+q_{2}\boldsymbol{\nu}}{p_{2}+q_{2}}(1+o(1))\quad\text{for}\,i\in\mathcal{C}_{a}\cap\mathcal{H}_{e},\\ -\frac{q_{2}\boldsymbol{\mu}+p_{2}\boldsymbol{\nu}}{p_{2}+q_{2}}(1+o(1))\quad\text{for}\,i\in\mathcal{C}_{n}\cap\mathcal{H}_{e}.\\ \end{aligned}\right.$ (21)
Applying the same linear classifier with $\boldsymbol{w}^{*}$ and $b^{*}$ yields the following results:
$\langle\boldsymbol{\hat{X}}_{i},\boldsymbol{w}^{*}\rangle+b^{*}=\left\{\begin{aligned} \frac{R(p_{2}-q_{2})}{2(p_{2}+q_{2})}\|\boldsymbol{\mu}-\boldsymbol{\nu}\|(1+o(1))<0\quad\text{for}\,i\in\mathcal{C}_{a}\cap\mathcal{H}_{e},\\ \frac{R(q_{2}-p_{2})}{2(p_{2}+q_{2})}\|\boldsymbol{\mu}-\boldsymbol{\nu}\|(1+o(1))<0\quad\text{for}\,i\in\mathcal{C}_{n}\cap\mathcal{H}_{e}.\\ \end{aligned}\right.$ (22)
Therefore, the same linear classifier is capable of separating both the homophilic node set $\mathcal{H}_{o}$ and the heterophilic node set $\mathcal{H}_{e}$. And according to part 2 of Theorem 1 in [2], we have
$\mathbb{P}\left(\left(\boldsymbol{\hat{X}_{i}}\right)_{i\in\mathcal{V}^{L}}\text{is linearly separable}\right)=1-o_{d}(1),$ (23)
where $o_{d}(1)$ denotes a quantity that converges to 0 as $d\to\infty$. ∎
Appendix E Formulations of Fusion Methods
@@
  • The "Mean" method averages the representations from the low-pass and high-pass encoders, i.e., $\boldsymbol{Z}=0.5\cdot(\boldsymbol{Z}_{L}+\boldsymbol{Z}_{H})$.
  • The "Concat" method concatenates the representations from the low-pass and high-pass encoders, i.e., $\boldsymbol{Z}=[\boldsymbol{Z}_{L},\boldsymbol{Z}_{H}]$.
  • The "Atten." method [6] employs an attention mechanism to learn the weights $\boldsymbol{\alpha}_{L},\boldsymbol{\alpha}_{H}\in[0,1]^{n\times 1}$ for $n$ nodes, such that $\boldsymbol{Z}=\boldsymbol{\alpha}_{L}\cdot\boldsymbol{Z}_{L}+\boldsymbol{\alpha}_{H}\cdot\boldsymbol{Z}_{H}$. Specifically, for node $v_{i}$ with $\boldsymbol{Z}_{i}^{L},\boldsymbol{Z}_{i}^{H}\in\mathbb{R}^{1\times e}$, the attention scores are computed as:
@@
Appendix F Time Complexity
@@
We first extract a Rayleigh Quotient-guided sub-tree $\mathcal{G}_{i}^{RQ}$ for each node using the MRQSampler [38], which has a total complexity of $O(n\log n)$. Since the sampling for each node is independent, this step can be parallelized to further reduce runtime. Importantly, this subgraph sampling is performed only once and reused in both training and inference, thereby introducing minimal overhead.
@@
In summary, the pre-training and fine-tuning phases of APF scale linearly with the graph size. Given that the sub-tree extraction is only performed once and can be computed in parallel, APF exhibits strong scalability and is well-suited for large-scale graphs. For example, our model can be applied to datasets like T-Social, which contains over 5.7 million nodes and 73 million edges. We additionally conduct efficiency comparison in Appendix G.5.
Appendix G Additional Experimental Details
G.1 Datasets
@@
Following GADBench [59], we conduct experiments on 10 real-world datasets spanning various scales and domains. Reddit [35], Weibo [35], Questions [51], and T-Social [60] focus on detecting anomalous accounts on social media platforms. Tolokers [51], Amazon [49], and YelpChi [55] are designed to identify fraudulent workers, reviews, and reviewers in crowdsourcing or e-commerce platforms. T-Finance [60], Elliptic [68], and DGraph-Fin [30] target the detection of fraudulent users, illicit entities, and overdue loans in financial networks. Dataset Statistics are presented in Table 3. Detailed descriptions of these datasets are as follows.
@@
  • Reddit [35]: This dataset includes a user-subreddit graph, capturing one month’s worth of posts shared across various subreddits. It includes verified labels for banned users and focuses on the 1,000 most active subreddits and the 10,000 most engaged users, resulting in 672,447 interactions. Posts are represented as feature vectors based on Linguistic Inquiry and Word Count (LIWC) categories.
@@
  • Weibo [35]: This dataset consists of a user-hashtag graph from the Tencent-Weibo platform, containing 8,405 users and 61,964 hashtags. Suspicious activities are defined as posting two messages within a specific time frame, such as 60 seconds. Users engaged in at least five such activities are labeled as "suspicious," while others are categorized as "benign." The feature vectors are based on the location of posts and bag-of-words features.
@@
  • Amazon [49]: This dataset focuses on detecting users who are paid to write fake reviews for products in the Musical Instrument category on Amazon.com. It contains three types of relationships: U-P-U (users reviewing the same product), U-S-U (users giving the same star rating within one week), and U-V-U (users with top 5% mutual review similarities).
@@
  • YelpChi [55]: This dataset aims to identify anomalous reviews on Yelp.com that unfairly promote or demote products or businesses. It includes three types of edges: R-U-R (reviews by the same user), R-S-R (reviews for the same product with the same star rating), and R-T-R (reviews for the same product within the same month).
@@
  • T-Finance [60]: This dataset is designed to detect anomalous accounts in transaction networks. The nodes represent unique anonymized accounts with 10-dimensional features related to registration days, login activities, and interaction frequency. Edges represent transactions between accounts, and anomalies are annotated by human experts based on categories such as fraud, money laundering, and online gambling.
@@
  • Elliptic [68]: This dataset includes over 200,000 Bitcoin transactions (nodes), 234,000 directed payment flows (edges), and 166 node features. It maps Bitcoin transactions to real-world entities, categorizing them as either licit (e.g., exchanges, wallet providers, miners) or illicit (e.g., scams, malware, terrorist organizations, ransomware, and Ponzi schemes).
@@
  • Tolokers [51]: This dataset is derived from the Toloka crowdsourcing platform. Nodes represent workers who have participated in at least one of 13 selected projects. An edge connects two workers if they collaborate on the same task. The goal is to predict which workers were banned in any of the projects. Node features are based on worker profiles and task performance.
@@
  • Questions [51]: This dataset is collected from the Yandex Q question-answering platform. It includes users as nodes, with edges representing answers between users during a one-year period (September 2021 to August 2022). It focuses on users interested in the "medicine" topic. The task is to predict which users remained active by the end of the period. Node features include the mean of FastText embeddings for words in the user descriptions, with a binary feature indicating users without descriptions.
@@
  • DGraph-Fin [30]: This dataset is a large-scale dynamic graph from the Finvolution Group representing a financial industry social network. Nodes represent users, and edges indicate emergency contact relationships. Anomalous nodes correspond to users exhibiting overdue behaviors. The dataset includes over 3 million nodes, 4 million dynamic edges, and more than 1 million unbalanced ground-truth anomalies.
@@
  • T-Social [60]: This dataset targets anomalous accounts in social networks. Nodes share the same annotations and features as those in T-Finance, with edges representing friend relationships maintained for more than three months. Anomalous nodes are annotated by experts in categories like fraud, money laundering, and online gambling.
@@
Table 3: Dataset statistics including the number of nodes and edges, the node feature dimension, the ratio of anomalies, the local homophily, the concept of relations, and the type of node features. $h^{a}$, and $h^{n}$ represent the average local homophily for abnormal nodes and normal nodes respectively. As shown, $h^{a}$ is much smaller than $h^{n}$, highlighting the presence of class-level local homophily disparity. "Misc." refers to node features that are a combination of heterogeneous attributes, which may include categorical, numerical, and temporal information.
@@ G.2 Baselines
We compare our model with a series of baseline methods, which can be categorized into the following groups: (1) Standard GNN Architectures, including GCN [34], GIN [69], GAT [62], ACM [45], FAGCN [4], AdaGNN [18], and BernNet [25]; (2) GNNs Specialized for GAD, including GAS [37], DCI [67], PCGNN [40], AMNet [6], BWGNN [60], GHRN [21], ConsisGAD [8], SpaceGNN [17], and XGBGraph [59]; (3) Graph Pre-Training Methods, including DGI [63], GRACE [72], G-BT [3], GraphMAE [29], BGRL [61], SSGE [43], PolyGCL [7], and BWDGI which incorporates BWGNN and DGI. Detailed descriptions of these baselines are as follows.
G.2.1 Standard GNN Architectures

### v3

@@ Appendix E Proof of Theorem 1
$\big|\langle\hat{\boldsymbol{X}}_{i}-\mathbb{E}[\hat{\boldsymbol{X}}_{i}\mid z_{i},h_{i}],\boldsymbol{w}\rangle\big|=\mathcal{O}\!\left(\sqrt{\frac{\log n}{d\,n\,\kappa_{\mathrm{eff}}}}\right),$ (14)
matching the CSBM rate up to the effective factor $\kappa_{\mathrm{eff}}$.
Pattern-dependent means after node-adaptive filtering.
Condition on $z_{i}$ and $h_{i}$. A neighboring node $v_{j}$ belongs to class $a$ with probability proportional to $\theta_{j}\beta^{(h_{i})}_{z_{i},a}$ and to class $n$ with probability proportional to $\theta_{j}\beta^{(h_{i})}_{z_{i},n}$. By (iii) and bounded heterogeneity $\chi=\mathcal{O}(1)$, the neighbor-class proportions concentrate at
$\omega_{i,a}^{(h_{i})}=\frac{\pi_{a}\beta^{(h_{i})}_{z_{i},a}}{\pi_{a}\beta^{(h_{i})}_{z_{i},a}+\pi_{n}\beta^{(h_{i})}_{z_{i},n}},\quad\omega_{i,n}^{(h_{i})}=1-\omega_{i,a}^{(h_{i})}.$
Thus the low-pass mean is a degree-normalized mixture
$\mathbb{E}[\boldsymbol{S}\boldsymbol{X}]_{i}=\omega_{i,a}^{(h_{i})}\boldsymbol{\mu}+\omega_{i,n}^{(h_{i})}\boldsymbol{\nu}\ (1\pm o(1)),$
while the high-pass mean flips the sign: $\mathbb{E}[-\boldsymbol{S}\boldsymbol{X}]_{i}=-\mathbb{E}[\boldsymbol{S}\boldsymbol{X}]_{i}$. Enumerating cases gives, for $h_{i}=o$ (homophily):
$\mathbb{E}[\hat{\boldsymbol{X}}_{i}]=\begin{cases}\frac{\pi_{a}p_{1}\boldsymbol{\mu}+\pi_{n}q_{1}\boldsymbol{\nu}}{\pi_{a}p_{1}+\pi_{n}q_{1}}\ (1\pm o(1)),&z_{i}=a,\\[4.0pt] \frac{\pi_{a}q_{1}\boldsymbol{\mu}+\pi_{n}p_{1}\boldsymbol{\nu}}{\pi_{a}q_{1}+\pi_{n}p_{1}}\ (1\pm o(1)),&z_{i}=n,\end{cases}$
and for $h_{i}=e$ (heterophily) the same expressions with $(p_{1},q_{1})$ replaced by $(p_{2},q_{2})$, followed by an overall sign flip due to the high-pass ($\boldsymbol{S}$). Consequently, under the node-adaptive choice (low-pass on $\mathcal{H}_{o}$ and high-pass on $\mathcal{H}_{e}$), all class-wise means align in the same ordering along direction $\boldsymbol{\nu}-\boldsymbol{\mu}$:
$\langle\mathbb{E}[\hat{\boldsymbol{X}}_{i}\mid z_{i}=a],\boldsymbol{\nu}-\boldsymbol{\mu}\rangle<\langle\mathbb{E}[\hat{\boldsymbol{X}}_{i}\mid z_{i}=n],\boldsymbol{\nu}-\boldsymbol{\mu}\rangle,$
with a gap proportional to $\Delta_{h}:=\frac{|p_{h}-q_{h}|}{\pi_{a}p_{h}+\pi_{n}q_{h}+\pi_{a}q_{h}+\pi_{n}p_{h}}$ (here $h\in\{1,2\}$ indexes $(p_{1},q_{1})$ or $(p_{2},q_{2})$), hence lower bounded by a constant multiple of $\frac{|p_{h}-q_{h}|}{p_{h}+q_{h}}$ and independent of $\theta$ due to the random-walk normalization.
Prior-aware linear separator and margin.
Consider the linear classifier with direction $\boldsymbol{w}^{*}=R\frac{\boldsymbol{\nu}-\boldsymbol{\mu}}{\|\boldsymbol{\mu}-\boldsymbol{\nu}\|}$ and a bias that accounts for the class prior shift
$b^{*}=-\frac{\langle\boldsymbol{\mu}+\boldsymbol{\nu},\boldsymbol{w}^{*}\rangle}{2}\ +\ \tau_{\pi},\qquad\tau_{\pi}=R\,\frac{\log(\pi_{a}/\pi_{n})}{\|\boldsymbol{\mu}-\boldsymbol{\nu}\|}.$
The additive term $\tau_{\pi}$ is the standard LDA correction under unequal priors with spherical covariances (Hastie et al., 2009). For any $v_{i}\in\mathcal{C}_{a}$, combining Step 2 and Eq 14 gives
$\langle\hat{\boldsymbol{X}}_{i},\boldsymbol{w}^{*}\rangle+b^{*}$ $=\langle\mathbb{E}[\hat{\boldsymbol{X}}_{i}\mid z_{i}=a],\boldsymbol{w}^{*}\rangle-\frac{\langle\boldsymbol{\mu}+\boldsymbol{\nu},\boldsymbol{w}^{*}\rangle}{2}+\tau_{\pi}+\underbrace{\langle\hat{\boldsymbol{X}}_{i}-\mathbb{E}[\hat{\boldsymbol{X}}_{i}],\boldsymbol{w}^{*}\rangle}_{\text{fluctuation}}$
$=-\frac{R}{2}\,\Gamma_{h_{i}}\,\|\boldsymbol{\mu}-\boldsymbol{\nu}\|\ (1\pm o(1))\ +\ \tau_{\pi}\ +\ \mathcal{O}\!\Big(R\sqrt{\tfrac{\log n}{d\,n\,\kappa_{\mathrm{eff}}}}\Big),$
where $\Gamma_{h_{i}}>0$ depends on the pattern (homophily vs. heterophily) via the mixture coefficients in Step 2 (and is uniformly bounded away from 0 as $|p_{h}-q_{h}|>0$). For $i\in\mathcal{C}_{n}$ the sign of the leading term is positive. Hence, if the feature center distance satisfies
$\|\boldsymbol{\mu}-\boldsymbol{\nu}\|\ \geq\ C\,\frac{\log n}{\sqrt{d\,n\,\kappa_{\mathrm{eff}}}}$
for a sufficiently large constant $C>0$, the margin term dominates both the stochastic fluctuation $\mathcal{O}\!\big(R\sqrt{\tfrac{\log n}{d\,n\,\kappa_{\mathrm{eff}}}}\big)$ and the prior correction $\tau_{\pi}=\mathcal{O}\!\big(R\,|\log(\pi_{a}/\pi_{n})|/\|\boldsymbol{\mu}-\boldsymbol{\nu}\|\big)$ even when $\pi_{a}\ll\pi_{n}$. Therefore, according to part 2 of Theorem 1 in Baranwal et al. (2021), $\operatorname{sign}(\langle\hat{\boldsymbol{X}}_{i},\boldsymbol{w}^{*}\rangle+b^{*})$ equals the true label for all nodes with probability $1-o_{d}(1)$, by a union bound over $v_{i}\in\mathcal{V}$. This establishes linear separability of node-adaptively filtered features under degree correction and class imbalance, proving the theorem. ∎
Appendix F Formulations of Fusion Methods
@@
  • The ”Mean” method averages the representations from the low-pass and high-pass encoders, i.e., $\boldsymbol{Z}=0.5\cdot(\boldsymbol{Z}_{L}+\boldsymbol{Z}_{H})$.
  • The ”Concat” method concatenates the representations from the low-pass and high-pass encoders, i.e., $\boldsymbol{Z}=[\boldsymbol{Z}_{L},\boldsymbol{Z}_{H}]$.
  • The ”Atten.” method Chai et al. (2022) employs an attention mechanism to learn the weights $\boldsymbol{\alpha}_{L},\boldsymbol{\alpha}_{H}\in[0,1]^{n\times 1}$ for $n$ nodes, such that $\boldsymbol{Z}=\boldsymbol{\alpha}_{L}\cdot\boldsymbol{Z}_{L}+\boldsymbol{\alpha}_{H}\cdot\boldsymbol{Z}_{H}$. Specifically, for node $v_{i}$ with $\boldsymbol{Z}_{i}^{L},\boldsymbol{Z}_{i}^{H}\in\mathbb{R}^{1\times e}$, the attention scores are computed as:
@@
Appendix G Time Complexity
@@
We first extract a Rayleigh Quotient-guided subgraph $\mathcal{G}_{i}^{RQ}$ for each node using the MRQSampler Lin et al. (2024), which has a total complexity of $O(n\log n)$. Since the sampling for each node is independent, this step can be parallelized to further reduce runtime. Importantly, this subgraph sampling is performed only once and reused in both training and inference, thereby introducing minimal overhead.
@@
In summary, the pre-training and fine-tuning phases of APF scale linearly with the graph size. Given that the subgraph extraction is only performed once and can be computed in parallel, APF exhibits strong scalability and is well-suited for large-scale graphs. For example, our model can be applied to datasets like T-Social, which contains over 5.7 million nodes and 73 million edges. We additionally conduct efficiency comparison in Appendix I.3.
Appendix H Additional Experimental Details
H.1 Datasets
@@
Following GADBench Tang et al. (2023), we conduct experiments on 10 real-world datasets spanning various scales and domains. Reddit Kumar et al. (2019), Weibo Kumar et al. (2019), Questions Platonov et al. (2023), and T-Social Tang et al. (2022) focus on detecting anomalous accounts on social media platforms. Tolokers Platonov et al. (2023), Amazon McAuley and Leskovec (2013), and YelpChi Rayana and Akoglu (2015) are designed to identify fraudulent workers, reviews, and reviewers in crowdsourcing or e-commerce platforms. T-Finance Tang et al. (2022), Elliptic Weber et al. (2019), and DGraph-Fin Huang et al. (2022) target the detection of fraudulent users, illicit entities, and overdue loans in financial networks. Dataset Statistics are presented in Table 4. Detailed descriptions of these datasets are as follows.
@@
  • Reddit Kumar et al. (2019): This dataset includes a user-subreddit graph, capturing one month’s worth of posts shared across various subreddits. It includes verified labels for banned users and focuses on the 1,000 most active subreddits and the 10,000 most engaged users, resulting in 672,447 interactions. Posts are represented as feature vectors based on Linguistic Inquiry and Word Count (LIWC) categories.
@@
  • Weibo Kumar et al. (2019): This dataset consists of a user-hashtag graph from the Tencent-Weibo platform, containing 8,405 users and 61,964 hashtags. Suspicious activities are defined as posting two messages within a specific time frame, such as 60 seconds. Users engaged in at least five such activities are labeled as ”suspicious,” while others are categorized as ”benign.” The feature vectors are based on the location of posts and bag-of-words features.
@@
  • Amazon McAuley and Leskovec (2013): This dataset focuses on detecting users who are paid to write fake reviews for products in the Musical Instrument category on Amazon.com. It contains three types of relationships: U-P-U (users reviewing the same product), U-S-U (users giving the same star rating within one week), and U-V-U (users with top 5% mutual review similarities).
@@
  • YelpChi Rayana and Akoglu (2015): This dataset aims to identify anomalous reviews on Yelp.com that unfairly promote or demote products or businesses. It includes three types of edges: R-U-R (reviews by the same user), R-S-R (reviews for the same product with the same star rating), and R-T-R (reviews for the same product within the same month).
@@
  • T-Finance Tang et al. (2022): This dataset is designed to detect anomalous accounts in transaction networks. The nodes represent unique anonymized accounts with 10-dimensional features related to registration days, login activities, and interaction frequency. Edges represent transactions between accounts, and anomalies are annotated by human experts based on categories such as fraud, money laundering, and online gambling.
@@
  • Elliptic Weber et al. (2019): This dataset includes over 200,000 Bitcoin transactions (nodes), 234,000 directed payment flows (edges), and 166 node features. It maps Bitcoin transactions to real-world entities, categorizing them as either licit (e.g., exchanges, wallet providers, miners) or illicit (e.g., scams, malware, terrorist organizations, ransomware, and Ponzi schemes).
@@
  • Tolokers Platonov et al. (2023): This dataset is derived from the Toloka crowdsourcing platform. Nodes represent workers who have participated in at least one of 13 selected projects. An edge connects two workers if they collaborate on the same task. The goal is to predict which workers were banned in any of the projects. Node features are based on worker profiles and task performance.
@@
  • Questions Platonov et al. (2023): This dataset is collected from the Yandex Q question-answering platform. It includes users as nodes, with edges representing answers between users during a one-year period (September 2021 to August 2022). It focuses on users interested in the ”medicine” topic. The task is to predict which users remained active by the end of the period. Node features include the mean of FastText embeddings for words in the user descriptions, with a binary feature indicating users without descriptions.
@@
  • DGraph-Fin Huang et al. (2022): This dataset is a large-scale dynamic graph from the Finvolution Group representing a financial industry social network. Nodes represent users, and edges indicate emergency contact relationships. Anomalous nodes correspond to users exhibiting overdue behaviors. The dataset includes over 3 million nodes, 4 million dynamic edges, and more than 1 million unbalanced ground-truth anomalies.
@@
  • T-Social Tang et al. (2022): This dataset targets anomalous accounts in social networks. Nodes share the same annotations and features as those in T-Finance, with edges representing friend relationships maintained for more than three months. Anomalous nodes are annotated by experts in categories like fraud, money laundering, and online gambling.
@@
Table 4: Dataset statistics including the number of nodes and edges, the node feature dimension, the ratio of anomalies, the local homophily, the concept of relations, and the type of node features. $h^{a}$, and $h^{n}$ represent the average local homophily for abnormal nodes and normal nodes respectively. As shown, $h^{a}$ is much smaller than $h^{n}$, highlighting the presence of class-level local homophily disparity. ”Misc.” refers to node features that are a combination of heterogeneous attributes, which may include categorical, numerical, and temporal information.
@@ H.2 Baselines
We compare our model with a series of baseline methods, which can be categorized into the following groups: (1) Standard GNN Architectures, including GCN Kipf and Welling (2017), GIN Xu et al. (2019), GAT Veličković et al. (2018), ACM Luan et al. (2022), FAGCN Bo et al. (2021), AdaGNN Dong et al. (2021), and BernNet He et al. (2021); (2) GNNs Specialized for GAD, including GAS Li et al. (2019), DCI Wang et al. (2021), PCGNN Liu et al. (2021a), AMNet Chai et al. (2022), BWGNN Tang et al. (2022), GHRN Gao et al. (2023), ConsisGAD Chen et al. (2024b), SpaceGNN Dong et al. (2025), and XGBGraph Tang et al. (2023); (3) Graph Pre-Training Methods, including DGI Veličković et al. (2019), GRACE Zhu et al. (2020), G-BT Bielak et al. (2022), GraphMAE Hou et al. (2022), BGRL Thakoor et al. (2022), SSGE Liu et al. (2024c), PolyGCL Chen et al. (2024a), and BWDGI which incorporates BWGNN and DGI. Detailed descriptions of these baselines are as follows.
H.2.1 Standard GNN Architectures